OpenStudy (anonymous):

Find volume of the solid generated by rotating the plane region bounded by x2−4x+2=y and y=2 about the line y=2. I'm having troubles finding the radius so i can do the question

OpenStudy (anonymous):

that' x^2 - 4x + 2 = y

OpenStudy (anonymous):

The radius is the whole equation integrated

OpenStudy (anonymous):

 look at the radius with respect to the representative rectangle. axis or revolution is y = 2. The height of the function is given by y = x^2 - 4x + 2. So, what is R = ?